OpenStudy (anonymous):

solve the following system of equations algebraical and check all solutions y=x2+3x-4 and y-x=7

OpenStudy (anonymous):

y = x + 7 x + 7 = x^2 + 3x - 4 x^2 + 2x - 11 = 0 quad formula x = -1 - 2 Sqrt[3],-1 + 2 Sqrt[3]

OpenStudy (anonymous):

Thank you soo much now all i do is subsitute those for X to check right :]

OpenStudy (anonymous):

yup